WORK AND ENERGY

The recently patented
Answer #1

(a) Work done by frictionless track is zero since the normal force is always making an angle of 90 with respect to displacment and since cos90=0 we have work done=0

(b) Using conservation of energy

m+2mg(2R) 15 × 202 +2 × 15 × 10(16) k = 2700 N/m

(c)

Radial accleraeration is given as =

\\ a_r=\frac{v^2}{r} \\ a_r=\frac{20^2}{8} \\ a_r=50 \ m/s^2 Downward

In this propblem component of tangential acceleration is very small and hence can be neglected

(d) Net Force at top

\\ \sum F_{net}=N+mg=\frac{mv^2}{R} \Rightarrow \frac{15\times 20^2}{8} = 750 \ N

(e)

\\ \sum F_{net}=N+mg = \frac{mv^2}{R} \\ N=\frac{mv^2}{R}-mg \\ N=750-150 \\ N=600 N

(f)

By conservation of energy

case 1

\\ 0.5kx^2=0.5mv^2 \\ v=x\sqrt{\frac{k}{m}}

case 2 x=2x

\\ 0.5k(2x)^2=0.5mv_n^2 \\ 4kx^2=mv_n^2 \\ v_n=2x\sqrt{\frac{k}{m}} \\ v_n=2v

So yeah on compressing the spring by twice more than before final impact velocity is doubled.